OpenStudy (anonymous):

What is the quotient of three and one over ten divided by two over five in simplest form? seven and three over four seven and one over two three and one over four three and one over twenty-five

OpenStudy (anonymous):

well you need to make it a multiplication problem 1 st

OpenStudy (ash2326):

We have \[3 \frac 1 {10}\ \ \div \frac 25\] could you simplify the dividend?

OpenStudy (ash2326):

@Pallavi06 Please don't just give answers. Explain the methodology you have used.

OpenStudy (anonymous):

\[3\frac{ 1 }{ 10 }= \frac{ 31 }{ 10} as (\frac{ (3\times10) +1 }{ 10 })\] now try....